# Public-private funding models in open source software development: A case study on scikit-learn

### v4

@@ Abstract
Governments are increasingly funding open source software (OSS) development to support software security, digital sovereignty, and national competitiveness in science and innovation, amongst others. However, little is known about how OSS developers evaluate the relative benefits and drawbacks of emergent governmental funding for OSS. This paper explores this question through a case study on scikit-learn, a popular Python library for machine learning, which has been funded by public research grants, commercial sponsorship, micro-donations, and a €32 million grant announced in France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, it contributes novel empirical findings on the effective design and implementation of a public-private funding model in an OSS project, as well as how the maintainers of scikit-learn have designed and employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects and makes recommendations to practitioners. The paper concludes with a discussion of the key recommendations.
@@ 1 Introduction
Open source software (OSS) has moved up the priority list of governments in light of concerns about digital sovereignty [1], software security [2], and national competitiveness in science and innovation [3], amongst others. The disclosure of the log4j vulnerability in November 2021 is widely seen as a turning point that mobilised national governments, from the United States of America (USA) to Germany, to mobilise public funds for the development and maintenance of OSS [4]. Increasing governmental involvement in funding OSS has been generally well-received by OSS developers. However, little is known about how they view the relative benefits and drawbacks of governmental funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use [5]. Its sustainability as a community-led project in the industry-dominated field of artificial intelligence (AI) [6] is partly attributed to its mixed funding model, including public research grants, commercial sponsorship, community donations, and the institutional backing from the French Institute for Research in Computer Science and Automation (Inria). In November 2021, in its national AI strategy, la “Stratégie IA”, the French government announced a €32 million grant to support the expansion of scikit-learn [7], marking a significant expansion of public funding for scikit-learn.
Through 25 interviews with scikit-learn’s maintainers and funders, this paper investigates the historical role of public and private funding in supporting the community-led OSS project. It explores the diverse interests of scikit-learn’s funders, and how they have aligned or conflicted with those of the maintainers (RQ1). The study also examines how the maintainers view the relative benefits and drawbacks of public and private funding (RQ2). Finally, it discusses practical lessons from scikit-learn’s funding model for OSS practitioners (RQ3). This paper makes two key contributions to research and practice. First, it provides novel empirical findings on the design and implementation of a public-private funding model in an OSS project. Furthermore, it sheds light on how the maintainers employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects. The paper concludes with a discussion of key recommendations to OSS developer communities, companies, and governments.
@@ 2.2 Private Funding for OSS Development
Recently, companies have begun to establish FOSS Contributor Funds, which allow employees to nominate and vote on OSS projects that should receive funding [41], and represent a more “democratic” way for companies to fund the OSS projects that they use [42]. A shortcoming is that these funds typically award up to \$10,000 over a one-year time frame. Alyssa Wright from Bloomberg’s FOSS Contributor Fund [43] has acknowledged this shortcoming, arguing that “true sustainability comes from a multi-sided commitment” and “while financial support is important, real engagement through contribution and dialogue is as crucial to the ecosystem’s survival as writing a check”.
@@ 2.4 Public Funding for OSS Development
The public sector has indirectly or directly funded OSS since the origins of the World Wide Web [65] through research grants for (scientific) OSS [66, 67, 68] as well as bespoke OSS funding bodies, such as the Open Technology Fund in the USA, the Next Generation Initiative at the European Commission, or the Sovereign Tech Fund (STF) in Germany [4]. Furthermore, OSS has recently moved up the priority list of governments in light of concerns about the security of OSS [69, 2]; its role as critical digital infrastructure [10, 11, 12]; and its importance for digital sovereignty [1, 70] and national competitiveness in science and innovation [3]. In Europe, the concern for digital sovereignty is particularly salient. For example, the European Commission’s OSS strategy argues that OSS can give “Europe a chance to create and maintain its own, independent digital approach and stay in control of its processes” [71].
Public sector support for OSS is also a recognised strategy to stimulate competition in domestic software markets, which might otherwise be monopolised by industry giants [72]. For example, it has been estimated that the Circulaire 5608, a French law requiring government agencies to favour OSS when procuring software, led to an increase of nearly 600,000 OSS contributions from French developers per year as well as yearly increases in the national competitiveness of the French IT market [3]. Beyond governmental adoption of OSS, government investments in OSS are often targeted at stimulating national competitiveness. For example, in June 2023, President Emmanuel Macron of France announced a €40 million fund to create a “digital commons” to support the development of open source large language models to support domestic startups and to challenge the dominance of Big Technology companies in the AI industry [73].
@@ 3.1 Research aims
This paper investigates the emerging role that governments are playing in funding OSS development and maintenance, and how OSS developers view the relative merits and drawbacks of governmental funding through a case study on the scikit-learn project. In software engineering research, a case study “is an empirical enquiry that draws on multiple sources of evidence to investigate one instance (or a small number of instances) of a contemporary software engineering phenomenon within its real-life context” [77]. Single case studies are suitable when one seeks to provide in-depth insights, generate new hypotheses, or develop theories about a critical, extreme, unusual, or revelatory phenomenon [78]. The scikit-learn project was selected as an exploratory case that could provide in-depth insights into the role of the French government in supporting this OSS project and how its maintainers view the relative merits and drawbacks of the government’s funding compared to the other funding sources that have supported the project to date.
@@ 3.2 Case presentation
scikit-learn is a Python library that implements ML algorithms for classification, regression, and clustering, as well as tools for data pre-processing, model evaluations, and data visualisation. Initially it was started as scikits.learn by David Courapeau during a Google Summer of Code project in 2007. After a dormant period, it was relaunched as scikit-learn by Fabian Pedregosa, Gaël Varoquaux, Alexandre Gramfort, and Vincent Michel with the backing of INRIA, with the first public release in February 2010. It has since become one of the most popular OSS libraries for ML [INRIA_2019_2020], described as “the Swiss army knife of ML” due to its widespread use [INRIA_2019_2020]. For example, in November 2023, it had 2,752 contributors and 61,085 GitHub stars [79]. The success and sustainability of this community-led OSS project is noteworthy, considering the extent of industry dominance in ML/AI research [6] and OSS ecosystems [80, 81].
scikit-learn’s sustainability is partly credited to its funding model, which has combined research grants, private sponsorships via the scikit-learn consortium, and community donations [82, 83]. The consortium was established under the Inria Foundation in 2018 in order to secure employment for maintainers, to sustain the quality of software, and to develop new features [83]. It comprises companies that pay annual membership fees. Since its founding, it has included Dataiku, Microsoft, Nvidia, Intel, AXA, Boston Consulting Group, BNP Paribas Cardif, Hugging Face, Nvidia, Fujitsu, and Chanel. Other companies, such as Quansight Labs, have indirectly funded the project by sponsoring maintainers. The project also benefits from donations from community members via NUMFOCUS and student projects have been sponsored by the Google Summer of Code programme. In November 2021, the French AI strategy, la “Stratégie IA”, announced a €32 million grant to support scikit-learn’s development [7]. This significant public investment alongside scikit-learn’s existing funding sources makes it a suitable case for the research aims of this exploratory case study.
@@ 4 Results
The results are organised into three parts, sequentially responding to RQ1 on how the interests of scikit-learn have aligned with or differed from those of the maintainers, and RQ2 on the relative benefits and drawbacks of public and private funding from the perspective of the scikit-learn maintainers. Finally, practical lessons and recommendations for practitioners based on the funding model of scikit-learn RQ3 are discussed in the subsequent Discussion section.
@@ 4.1 Stewards of the Community: The Roles & Interests of the scikit-learn Maintainers
scikit-learn’s maintainers are a diverse group employed by INRIA, public universities, and software companies. While some members of the core team work on scikit-learn full-time, others balance their roles as maintainers with their full-time jobs. Several maintainers also held academic positions at universities and research labs. The maintainers explained they perform a range of technical and social tasks, including reviewing code, mentoring contributors, and community building. François Goupil, the community manager, explained, “Of course, software development itself is the major part of the workload—going through issues, pull requests, and responding to community feedback—but lots of other work goes into the project. Community animation is a huge part. After all, it’s a community project”. In addition, sponsored developers from companies make focused contributions, such as the development of new features, which are agreed upon by the maintainers and the sponsoring companies. For example, at the time of this research, Nvidia sponsord a full-time maintainer who was building a plug-in to enable alternative computational backends for scikit-learn to enable compatibility with Nvidia’s GPU AI hardware accelerators.
Several maintainers expressed a strong sense of stewardship for the project, whose strength they viewed as rooted in its diverse community, comprising contributors from across the world (see Table 1) and companies (see Table 2). Julien Jerphanion, a maintainer, highlighted this community ethos, noting that, “the major part of the labour [in the project] is based on benevolence of people working in their free time and not asking to get paid”. Furthermore, Gaël Varoquaux, a co-founder and maintainer, explained that the community was at the heart of the project’s DNA and that major decisions about the future of the project remained community-driven despite their diverse funders. Similarly, Adrin Jalali, a maintainer, commented, “Even as a [maintainer], I don’t have enough credit to just be like, ‘Hey, we should do this.’ It’s a slow project with hundreds of 1000s of users. You can’t just change things”.
@@
Table 2: Most active countries in the scikit-learn community [79]
@@
In addition, the maintainers expressed their personal interests in maintaining scikit-learn. Several maintainers explained they became maintainers because of their scientific interest in ML, and they liked the scientific culture of the project. Many of the maintainers had academic backgrounds, some still worked in research labs, and some still taught at universities. For example, Julien Jerphanion said that he enjoyed working at the intersection of statistical theory and software engineering, which was lacking in his previous job in industry. In addition, an anonymous maintainer explained that while he may not become a millionaire, he had a comfortable life and enjoyed the challenge of the job. For many, being a maintainer of scikit-learn was also a source of pride, given the project’s reputation.
Beyond maintaining scikit-learn, several maintainers mentioned the importance for them to contribute to the wider scientific Python ecosystem and to support their dependencies, such as numpy and scipy. During my on-site visit in September 2022, the maintainers had just returned from EuroSciPy, an annual gathering of the European scientific Python community. One anonymous maintainer spoke fondly of “putting faces to GitHub IDs”. François Goupil explained that beyond collaborating on the code base, they collaborate with other projects in other ways, such as by applying for grants together. In addition, some of the maintainers had initiated “satellite projects”, such as joblib and skrub. However, their ability to support the ecosystem is constrained by their limited capacity. Julien Jerphanion explained that while projects like scikit-learn, numpy, and scipy are very important in ML/AI, “There are many which are not visible to the public, for example OpenBLAS which is fundamental and only maintained by two people, who have very specific expertise. This is tremendous software work, but it is invisible to the public. Projects like this are cornerstones for many projects. They are the foundations; they are, like, infrastructure — you do not see them if there is no problem in your daily life”.
4.2 Public Funding: From Research Grants at INRIA to “la Stratégie IA”
@@
Public funding has been a cornerstone of scikit-learn’s development since its beginning. Starting as a Google Summer of Code project in 2007, the library was later revived by researchers at INRIA. Over the past 13 years, INRIA had provided an estimated €1.5 million in support through staff salaries, research grants, office space, computing resources, and event sponsorship. As Olivier Grisel, a maintainer, noted, “[Public funding] is not new with the AI strategy”. However, the announcement of a €32 million grant in France’s AI strategy, la “Stratégie IA”, supporting the expansion of scikit-learn and OSS tools for data science, marked a significant expansion of public funding for scikit-learn. A government official explained that the grant seeks to facilitate AI adoption across the economy and to enhance French competitiveness in AI R&D, while also supporting the “digital sovereignty” of France and Europe.
@@
The maintainers were delighted to receive the funding and recognition from the government. Adrin Jalali highlighted the stability it provided for the project, enabling longer-term planning around hiring and the technical roadmap. Julien Jerphanion characterised it as “a good signal that the French state is getting involved in funding projects like this”. François Goupil particularly commended the requirement to seek matching funds from both public and private sources across the European Union, viewing this as an important safeguard for the project’s independence: “I think it would be dangerous for us to be exclusively funded by the private sector or to be exclusively funded by the French government, because we have many good contributors who are not French”.
@@
At the same time, challenges have emerged in aligning broad policy goals with the day-to-day realities and expertise of the maintainers. Initial requests from officials to expand scikit-learn into deep learning to provide a “sovereign” alternative to PyTorch and TensorFlow were resisted, with maintainers insisting on focusing on the project’s core strength in classical ML and the need to allocate sufficient funding and capacity to maintenance. The government officials ultimately deferred to their expertise, with one official acknowledging, “We understand that scikit-learn is specialised in ML, and it should not compete against TensorFlow and PyTorch”.
@@
However, other tensions have proven thornier. Requirements to simultaneously develop OSS tools for data science whilst not competing with domestic companies have left maintainers in a bind, as the resulting tools inherently provide open source alternatives to the proprietary offerings of companies. Maintainers have also expressed frustration with the slow pace of disbursement, with funds still not arriving 17 months after the announcement by the time of the last interview. While scikit-learn’s diverse funding model provided a buffer, they worried that such sluggishness could threaten the sustainability of more financially precarious OSS projects.
@@
Another official explained that the funding decision was informed by a multi-stakeholder process involving experts from research and industry. “It was not just the government. It was a dialogue between several actors—with the AI strategy team, the ‘Confiance IA’ project,111“Confiance IA” is a government-sponsored programme that aims to develop and industrialise trustworthy AI technologies in France. INRIA, with scikit-learn, and so on”. He confessed they had faced hurdles within the government: “Financing open source means financing a product that cannot be valued on the market and a priori it is not clear what its success will be. So, if [we want] the government to finance it in a substantial way—here we are talking about €32 million—the government wants to know where the money is going and to achieve a measurable return on investment”. Ultimately, it took months of coordination and negotiation to get the grant approved by the finance ministry.
@@
Two consortium members raised concerns that such a substantial government grant could politicise scikit-learn, potentially influencing it to prioritise French policy objectives over the interests of the community. One anonymous representative underlined the need to balance the French government’s concern about digital sovereignty with the norms of the global community, while another argued that imposing French policy priorities on the project would not be received well by the community. These concerns were acknowledged by the maintainers, who insisted that the community ethos of the project remained at the heart of their discussions with the government.

### v5

@@ Abstract
Governments are increasingly funding open source software (OSS) development to support software security, digital sovereignty, and national competitiveness in science and innovation, amongst others. However, little is known about how OSS developers evaluate the relative benefits and drawbacks of governmental funding for OSS. This study explores this question through a case study on scikit-learn, a Python library for machine learning, funded by public research grants, commercial sponsorship, micro-donations, and a €32 million grant announced in France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions. First, it contributes empirical findings about the benefits and drawbacks of public and private funding in an impactful OSS project, and the governance protocols employed by the maintainers to balance the diverse interests of their community and funders. Second, it offers practical lessons on funding for OSS developers, governments, and companies based on the experience of scikit-learn. The paper concludes with key recommendations for practitioners and future research directions.
@@ 1 Introduction
Open source software (OSS) has moved up the priority list of governments in light of concerns about digital sovereignty [1], software security [2], and national competitiveness in science and innovation [3], amongst others. The discovery of the log4j vulnerability in November 2021 was a turning point that animated discussion about OSS funding. From the White House’s software security meeting in January 2022 [74] to the establishment of the Sovereign Tech Fund (STF) in Germany in October 2022 [76], there has been a growing recognition amongst policymakers of the public value of funding the development and maintenance of OSS, increasingly recognised as digital infrastructure underpinning our digital society and economy [4, 10]. Increasing governmental interest and involvement in funding OSS has been generally well-received by OSS developers. However, little is known about how they view the relative benefits and drawbacks of governmental funding compared to other funding sources that have funded OSS developers to date.
This study explores this question through a case study on scikit-learn, a Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use in research and industry [5]. Its sustainability as a community-led project in the industry-dominated field of artificial intelligence (AI) [6] is partly attributed to its mixed funding model, including public research grants, commercial sponsorship, community donations, and the institutional backing from the French Institute for Research in Computer Science and Automation (Inria). In November 2021, in its national AI strategy, la Stratégie IA, the French government announced a €32 million grant to support the expansion of scikit-learn [7], marking a significant expansion of public funding for scikit-learn.
Through 25 interviews with scikit-learn’s maintainers and funders, this study investigates the historical role of public and private funding in supporting the community-led OSS project. This study makes two key contributions to research and practice. First, it contributes empirical findings about the benefits and drawbacks of public and private funding in an OSS project. Furthermore, it sheds light on how the maintainers employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects for OSS developers, companies, and governments.
@@ 2.2 Private Funding for OSS Development
Recently, companies have begun to establish FOSS Contributor Funds, which allow employees to nominate and vote on OSS projects that should receive funding [41], and represent a more democratic way for companies to fund the OSS projects that they use [42]. A shortcoming is that these funds typically award up to \$10,000 over a one-year time frame. Alyssa Wright from Bloomberg’s FOSS Contributor Fund [43] has acknowledged this shortcoming, arguing that “true sustainability comes from a multi-sided commitment” and “while financial support is important, real engagement through contribution and dialogue is as crucial to the ecosystem’s survival as writing a check.”
@@ 2.4 Public Funding for OSS Development
The public sector has indirectly or directly funded OSS since the origins of the World Wide Web [65] through research grants for (scientific) OSS [66, 67, 68] as well as bespoke OSS funding bodies, such as the Open Technology Fund in the United States of America (USA), the Next Generation Initiative at the European Commission, or the STF in Germany [4]. Furthermore, OSS has recently moved up the priority list of governments in light of concerns about the security of OSS [69, 2]; its role as critical digital infrastructure [10, 11, 12]; and its importance for digital sovereignty [1, 70] and national competitiveness in science and innovation [3]. In Europe, the concern for digital sovereignty is particularly salient. For example, the European Commission’s OSS strategy argues that OSS can give “Europe a chance to create and maintain its own, independent digital approach and stay in control of its processes” [71].
Public sector support for OSS is also a recognised strategy to stimulate competition in domestic software markets, which might otherwise be monopolised by industry giants [72]. For example, it has been estimated that the Circulaire 5608, a French law requiring government agencies to favour OSS when procuring software, led to an increase of nearly 600,000 OSS contributions from French developers per year as well as yearly increases in the national competitiveness of the French IT market [3]. Beyond governmental adoption of OSS, government investments in OSS are often targeted at stimulating national competitiveness. For example, in June 2023, President Emmanuel Macron of France announced a €40 million fund to create a digital commons to support the development of open source large language models to support domestic startups and to challenge the dominance of Big Technology companies in the AI industry [73].
@@ 3.1 Research Aims
This study investigates the emerging role that governments are playing in funding OSS development and maintenance, and how OSS developers view the relative merits and drawbacks of governmental funding through a case study on the scikit-learn project. In software engineering research, a case study “is an empirical enquiry that draws on multiple sources of evidence to investigate one instance (or a small number of instances) of a contemporary software engineering phenomenon within its real-life context” [77]. Single case studies are suitable when one seeks to provide in-depth insights, generate new hypotheses, or develop theories about a critical, extreme, unusual, or revelatory phenomenon [78]. The scikit-learn project was selected as an exploratory case that could provide in-depth insights into the role of the French government in supporting this OSS project and how its maintainers view the relative merits and drawbacks of the government’s funding compared to the other funding sources that have supported the project to date.
@@ 3.2 Case Presentation
scikit-learn is a Python library that implements ML algorithms for classification, regression, and clustering, as well as tools for data pre-processing, model evaluations, and data visualisation. Initially it was started as scikits.learn by David Courapeau during a Google Summer of Code project in 2007. After a dormant period, it was relaunched as scikit-learn by Fabian Pedregosa, Gaël Varoquaux, Alexandre Gramfort, and Vincent Michel with the backing of Inria, with the first public release in February 2010. It has since become one of the most popular OSS libraries for ML, described as “the Swiss army knife of ML” due to its widespread use [5]. For example, in November 2023, it had 2,752 contributors and 61,085 GitHub stars [79]. The success and sustainability of this community-led OSS project is noteworthy, considering the extent of industry dominance in ML/AI research [6] and OSS ecosystems [80, 81].
scikit-learn’s sustainability is partly credited to its mixed funding model [82, 83]. First, Inria has funded scikit-learn since 2010, including by paying salaries, administering research grants, and providing office space, amongst others. Second, community members have made micro-donations via NUMFOCUS and student projects have been sponsored by the Google Summer of Code programme. Third, the scikit-learn consortium, established under the Inria Foundation in 2018, offers three levels of annual memberships for companies: silver (€30,000), gold (€50,000), and platinum (€100,000), each providing varying degrees of involvement in the Technical and Advisory Committees. The Technical Committee develops the strategic technical roadmap for the project, which involves collecting input from the community, while the Advisory Committee provides guidance on strategic matters and the consortium’s membership, amongst others. Since its founding, it has included Dataiku, Microsoft, Nvidia, Intel, AXA, Boston Consulting Group, BNP Paribas Cardif, Hugging Face, Nvidia, Fujitsu, and Chanel. Companies, including Quansight Labs, Nvidia, and HuggingFace, have also sponsored maintainers. Fourth, in November 2021, the French AI strategy, la Stratégie IA, announced a €32 million grant to support scikit-learn [7]. This significant public investment, alongside other funding sources, makes scikit-learn a suitable case for the research aims of this case study.
@@ 4 Results
The results are organised into three parts, sequentially responding to RQ1 on how funders’ interests align with or differ from those of maintainers and RQ2 on the relative benefits and drawbacks of public and private funding from the perspective of the maintainers. Finally, practical lessons and recommendations for practitioners, as per RQ3, are discussed in the subsequent Discussion section.
@@ 4.1 Stewards of the Community: The Roles & Interests of the scikit-learn Maintainers
scikit-learn’s maintainers are a diverse group employed by Inria, public universities, and software companies. While some core team members work on scikit-learn full-time, others balance maintainer roles with their day jobs. Several maintainers also hold academic positions. The maintainers explained they perform a range of technical and social tasks, including reviewing code, mentoring contributors, and community building. François Goupil, the community manager, explained, “Of course, software development itself is the major part of the workload—going through issues, pull requests, and responding to community feedback—but lots of other work goes into the project. Community animation is a huge part.” In addition, company-sponsored developers make focused contributions, such as the development of new features, which are agreed upon by the maintainers and the respective companies. For example, Nvidia sponsors a full-time maintainer, who is building a plug-in to enable alternative computational backends for scikit-learn to enable compatibility with Nvidia’s GPU AI hardware accelerators.
Several maintainers expressed a strong sense of stewardship for the project’s diverse community, comprising contributors from across the world (see Table 1) and companies (see Table 2). Gaël Varoquaux, a co-founder, explained that the community was at the heart of the project’s DNA and underlined the importance of taking a community-driven approach to major decisions about the future of the project. Julien Jerphanion, a maintainer, highlighted the importance of community contributions, “The major part of the labour [in the project] is based on benevolence of people working in their free time and not asking to get paid.” Meanwhile Adrin Jalali, a maintainer, commented on the impact of micro-donations from community members, explaining that while “the major stuff is not funded through NUMFOCUS”, it has been useful in funding marketing, events, and an internship for underrepresented groups.
In addition, several maintainers expressed their scientific interest in ML and the project’s scientific culture as important incentives for contributing to scikit-learn. Many maintainers had academic backgrounds, worked in research labs, or taught at universities. For example, Jerphanion said that he enjoyed working at the intersection of statistical theory and software engineering, which was lacking in his previous industry job. In addition, an anonymous maintainer explained that while he may not become a millionaire, he had a comfortable life and enjoyed the job’s challenge. Furthermore, for many, being a scikit-learn maintainer was a source of pride, given scikit-learn’s global reputation and impact.
Beyond maintaining scikit-learn, several maintainers discussed the importance of contributing to their dependencies, such as numpy and scipy, and the wider scientific Python ecosystem. Goupil explained that they also collaborate with these projects by applying for grants together. In addition, some maintainers had initiated satellite projects, such as joblib and skrub. However, their ability to support the ecosystem is constrained by limited capacity. Jerphanion explained that while projects like scikit-learn, numpy, and scipy are very important in ML/AI, “There are many which are not visible to the public, for example OpenBLAS which is fundamental and only maintained by two people, who have very specific expertise. This is tremendous software work, but it is invisible to the public. Projects like this are cornerstones for many projects. They are the foundations; they are the infrastructure.”
@@
Table 2: Most Active Companies in the scikit-learn Community [79]
@@
4.2 Public Funding: From Research Grants at Inria to la Stratégie IA
@@
Public funding has supported the development of scikit-learn since its beginning in 2010. Over 13 years since its first public release, Inria has provided an estimated €1.5 million through staff salaries, research grants, office space, computing resources, and event sponsorship. This led Olivier Grisel, a maintainer, to explain that, “[Public funding] is not new with the AI strategy.” However, the €32 million grant in France’s AI strategy, la Stratégie IA, supporting the expansion of scikit-learn and the development of OSS tools for data science, marked a significant expansion of public funding. The deputy coordinator of the AI strategy explained that the grant seeks to facilitate AI adoption, enhance French competitiveness in AI R&D, and support the digital sovereignty of France and Europe.
@@
The maintainers were delighted to receive the recognition from the government. Jalali highlighted the stability it provided for the project, enabling long-term planning around hiring and the technical roadmap. Jerphanion saw it as “a good signal that the French state is getting involved in funding projects like this.” Goupil especially commended the requirement to seek matching funds from public and private sources across Europe, viewing it as an important safeguard for the project’s independence: “I think it would be dangerous for us to be exclusively funded by the private sector or to be exclusively funded by the French government, because we have many good contributors who are not French.”
@@
At the same time, challenges have emerged in aligning broad policy goals with the day-to-day realities and expertise of the maintainers. Initial requests from the government officials to expand scikit-learn into deep learning to provide a “sovereign” alternative to PyTorch and TensorFlow were resisted, with the maintainers insisting on focusing on the project’s specialism in ML and the need to allocate sufficient funding to maintenance. The coordinators of the AI strategy deferred to their expertise, with one official acknowledging, “We understand that scikit-learn is specialised in ML, and it should not compete against TensorFlow and PyTorch.”
@@
However, other demands have proven trickier. Requirements to simultaneously develop OSS tools for data science whilst not competing with domestic companies have left maintainers in a bind, as the resulting tools will inherently provide free alternatives to the proprietary offerings of companies. Maintainers have also expressed frustration with the slow pace of disbursement, with funds still not arriving 17 months after the announcement by the time of the last interview. While scikit-learn’s diverse funding model provided a buffer, they worried that such sluggishness could threaten the sustainability of more financially precarious OSS projects.
@@
One of the government officials explained that the decision to fund scikit-learn, as well as how to fund it, was informed by an ongoing multi-stakeholder process. “It was a dialogue between several actors—with the AI strategy team, the ‘Confiance IA’ project,111“Confiance IA” is a government-sponsored programme that aims to develop trustworthy AI technologies in France. Inria, with scikit-learn, and so on.” He confessed they had faced hurdles within the government: “Financing open source means financing a product that cannot be valued on the market and a priori it is not clear what its success will be. So, if [we want] the government to finance it in a substantial way—here we are talking about €32 million—the government wants to know where the money is going and to achieve a measurable return on investment.” Ultimately, it took months of coordination to get the grant approved by the finance ministry.
@@
Two consortium members raised concerns that such a substantial grant could politicise scikit-learn, potentially influencing it to prioritise French policy goals over the interests of the community. One representative underlined the need to balance the French government’s concern about digital sovereignty with the norms of the global community, while another argued that imposing French policy priorities on the project would not be received well by the community. These concerns were acknowledged by the maintainers, who insisted that the community ethos of the project remained at the heart of their discussions with the government officials.